# Designing quantum chemistry algorithms with Just-In-Time compilation

### v2

@@ 1. Introduction
Beyond runtime performance, just-in-time compilation also offers substantial advantages to developers. JIT facilitates rapid prototyping and testing: since code is generated at runtime, one can easily benchmark different kernel variants, tiling strategies, or instruction sequences, without recompiling and linking the entire application. This accelerates development cycles, simplifies debugging, and lowers the barrier to contributing performance improvements across architectures. Within a narrow area of application such as machine learning, SQL and GPU programming, one can design domain-specific languages (DSLs) via LLVM (LA, 04). Frameworks like TensorFlow(AAB+, 15), PyTorch(PGM+, 19), JAX(BFH+, 18), and TVM incorporate their own DSLs—such as TensorFlow’s GraphDef and XLA HLO, PyTorch’s TorchScript and FX, JAX’s JAXpr, and TVM’s TIR—to represent algorithms in intermediate forms that enable aggressive optimization and efficient execution. These DSLs allow developers to write high-level code while benefiting from low-level performance tuning through JIT compilation and backend-specific code generation, bridging the gap between productivity and performance.
JIT compilation benefits mixed precision algorithms in quantum chemistry, which naturally accelerate electron repulsion integrals by exploiting the Schwarz inequality. Most tasks can be calculated in single-precision (FP32) as compared to the commonly-used double-precision (FP64). Single precision algorithms benefit quantum-chemistry computations in three key ways: (1) deep learning inference GPUs and consumer-grade GPUs typically feature far more FP32 units than FP64 units, consume less power, and are more readily available; (2) single-precision data occupy half the storage, allowing more values to reside in shared memory and registers, improving data locality and throughput; and (3) hardware-accelerated FP32 implementations of the exponential (exp) and error function (erf) offer 4×–8× speedups, alleviating what is a computational bottleneck in certain workloads. JIT compilation unlocks these benefits without significant changes to the code.
@@ 2. A short formulation of Rys quadrature
where and are the roots and weights of Rys quadrature. The roots and weights depend on the exponents and center (atomic) positions of the four primitive Gaussian functions, the total angular momentum, and a range-separation parameter if needed. The calculation of Rys parameters relies on interpolations from tabulated data, and the details have been documented in (FL, 08). The primitive integral at each root position can be factored into intermediate variables in three spatial directions:
@@
, , and are the intermediate variables introduced in (FL, 08). Each of them is a tensor of shape . They depend on the exponents, atomic coordinates and angular momentum of the four primitive Gaussians only. They can be efficiently evaluated using recursive relations, documented in (FL, 08). But storing those values on the limited resources on GPU can be tricky, and several strategies will be discussed in Sec. 3.1 and Sec. 3.2. The index set for enumerating the Guassian functions of a primitive shell is defined as
@@ 3. Design and Algorithms
| Template Parameter | Data Type | Description |
| int | Angular momentum of shells | |
| do_j | bool | Whether to compute the J matrix |
| do_k | bool | Whether to compute the K matrix |
| rys_type | int | 0: no range separation (full range), |
| 1: long-range, -1: short-range | ||
| n_dm | int | # of density matrices |
| (e.g. two for unrestricted DFT) | ||
| DataType | type | float or double precision |
| int | # of primitives in the | |
| (i, j, k, l) shells | ||
| int | Fragment sizes of the | |
@@
| Input | Data Type | Description |
| double/single | contraction coefficients in basis | |
| double/single | exponents in basis | |
| coords | double/single | coordinates of atoms |
| double | density matrix as input | |
| double | coulomb potential as output | |
| double | exchange potential as output | |
| int | # of atomic orbitals | |
@@ 3.2. Quartet fragmentation (1qnt) algorithm
For high-angular momentum integrals, the intermediate variables and integral elements of one shell quartet cannot fit in the registers. From GPU4PySCF v1.0-v1.4, a fixed-size 1D fragmentation strategy is introduced to split the integral into multiple threads. Instead of one block thread processing all integrals of one shell quartet, a group of 2 to 256 threads cooperatively compute all integral elements within one shell quartet. The strategy significantly improved the performance of the high-angular integrals. In this work, we introduce the dynamic fragmentation strategy to further improve the data locality, and multi-level reduction algorithm to reduce the bottleneck of memory access (both shared memory and global memory). Since multiple threads for the same shell quartet use the same intermediate variables, the intermediate variable , and will be stored in shared memory. Then, each thread in the group will compute and store into registers only a fraction of the integrals. For example, the shape of the integral (dd—dd) is (6,6,6,6), each thread computes a fragment with the size of (3,3,3,3). 16 threads are needed to store the entire (dd—dd) integral. In the reduction stage, each thread is responsible for a fragment of 3x3 and . The fragments and are further reduced at the thread block level. The reduction scheme is presented in Algorithm 1. The shape of the intermediate variables are also known at compilation time.
@@ 4.1. Compilation Overhead
| Integral pattern | Algorithm | Compilation time (ms) | Binary size (KB) |
| (ss—ss)/(11—11) | 1Q1T | 284.8 | 25 |
| (ss—ss)/(33—33) | 1Q1T | 512.0 | 74 |
| (pp—pp)/(11—11) | 1Q1T | 638.8 | 75 |
| (pp—pp)/(33—33) | 1Q1T | 880.6 | 94 |
| (dd—dd)/(11—11) | 1QnT | 6011 | 194 |
| (ff—ff)/(11—11) | 1QnT | 7697 | 270 |
@@ 4.2. Benchmark with GPU4PySCF and TeraChem
| Molecule | Basis | GPU4PySCF/FP64 - xQC/FP32 | GPU4PySCF/FP64 - TeraChem/FP64 | xQC/FP32 - TeraChem/FP32 |
| gly30 | 6-31G* | 0.23 mHa | 0.0005 mHa | -0.07 mHa |
| gly30 | def2-TZVPP | 1.95 mHa | 0.0004 mHa | 2.10 mHa |
| valinomycin | 6-31G* | -0.06 mHa | -0.001 mHa | -0.26 mHa |
@@ Acknowledgements.
The authors gratefully acknowledge the support of Dr. Wen Yan and the ByteDance Seed management team. Special thanks are also extended to Dr. Qiming Sun and Dr. Zhichen Pu for their insightful discussions. Dr. Qiming contributed to the improvements of JK kernels v1.0 to v1.4 in GPU4PySCF. The integral algorithm before GPU4PySCF v1.0 is documented in (LSZC, 25). The authors further thanks the PySCF community for providing valuable tools that contributed to this work.References

### v4

@@ 1. Introduction
Beyond runtime performance, just-in-time compilation also offers substantial advantages to developers. JIT facilitates rapid prototyping and testing: since code is generated at runtime, one can easily benchmark different kernel variants, tiling strategies, or instruction sequences, without recompiling and linking the entire application. This accelerates development cycles, simplifies debugging, and lowers the barrier to contributing performance improvements across architectures. Furthermore, the developers are able to design novel algorithms which are not general enough under the AOT framework. In this paper, we demonstrate several algorithms due to this benefit. The development experience can be even further improved: Within a narrow area of application such as machine learning, SQL and GPU programming, one can design domain-specific languages (DSLs) via LLVM (LA, 04). Frameworks like TensorFlow(AAB+, 15), PyTorch(PGM+, 19), JAX(BFH+, 18), and TVM incorporate their own DSLs—such as TensorFlow’s GraphDef and XLA HLO, PyTorch’s TorchScript and FX, JAX’s JAXpr, and TVM’s TIR—to represent algorithms in intermediate forms that enable aggressive optimization and efficient execution. These DSLs allow developers to write high-level code while benefiting from low-level performance tuning through JIT compilation and backend-specific code generation, bridging the gap between productivity and performance.
JIT compilation benefits mixed precision algorithms in quantum chemistry, which naturally accelerate electron repulsion integrals by exploiting the Schwarz inequality (LUM, 11; MM, 13). Most tasks can be calculated in single-precision (FP32) as compared to the commonly-used double-precision (FP64). Single precision algorithms benefit quantum-chemistry computations in three key ways: (1) deep learning inference GPUs and consumer-grade GPUs typically feature far more FP32 units than FP64 units, consume less power, and are more readily available; (2) single-precision data occupy half the storage, allowing more values to reside in shared memory and registers, improving data locality and throughput; and (3) hardware-accelerated FP32 implementations of the exponential (exp) and error function (erf) offer 4×–8× speedups, alleviating what is a computational bottleneck in certain workloads. These benefits can be better utilized by JIT compilation without significant changes to the code.
@@ 2. A short formulation of Rys quadrature
where is the number of Rys roots. and are the roots and weights of Rys quadrature. The roots and weights depend on the exponents and center (atomic) positions of the four primitive Gaussian functions, the total angular momentum, and a range-separation parameter if needed. The calculation of Rys roots and weights relies on interpolations from tabulated data, and the details have been documented in (FL, 08). The number of Rys roots is determined at the compilation time, and only the data(WSP+, 25; Sun, 15) associated with the number of roots are compiled into the binary. The primitive integral at each root position can be factored into intermediate variables in three spatial directions:
@@
, , and are the intermediate variables introduced in (DRK, 76; FL, 08). Each of them is a tensor of shape . They depend on the exponents, atomic coordinates and angular momentum of the four primitive Gaussians only. They can be efficiently evaluated using recursive relations, documented in (FL, 08). But storing those values on the limited resources on GPU can be tricky, and several strategies will be discussed in Sec. 3.1 and Sec. 3.2. The index set for enumerating the Guassian functions of a primitive shell is defined as
@@ 3. Design and Algorithms
Data Type | Description | ||
| int | Angular momentum of shells | |||
| do_j | bool | Whether to compute the J matrix | ||
| do_k | bool | Whether to compute the K matrix | ||
| rys_type | int | 0: no range separation (full range), | ||
| 1: long-range, -1: short-range | ||||
| n_dm | int | # of density matrices | ||
| (e.g. two for unrestricted DFT) | ||||
| DataType | type | float or double precision | ||
| int | # of primitives in the | |||
| (i, j, k, l) shells | ||||
| int | Fragment sizes of the | |||
@@
Data Type | Description | |
| double/single | contraction coefficients in basis | ||
| double/single | exponents in basis | ||
| coords | double/single | coordinates of atoms | |
| double | density matrix as input | ||
| double | coulomb potential as output | ||
| double | exchange potential as output | ||
| int | # of atomic orbitals | ||
@@ 3.2. Quartet fragmentation (1qnt) algorithm
For high-angular momentum integrals, the intermediate variables and integral elements of one shell quartet cannot fit in the registers. From GPU4PySCF v1.0-v1.4, a fixed-size 1D fragmentation strategy is introduced to split the integral into multiple threads. Instead of one block thread processing all integrals of one shell quartet, a group of 2 to 256 threads cooperatively compute all integral elements within one shell quartet. The strategy significantly improved the performance of the high-angular integrals. In this work, we introduce the dynamic fragmentation strategy to further improve the data locality, and multi-level reduction algorithm to reduce the bottleneck of memory access (both shared memory and global memory). Since multiple threads for the same shell quartet use the same intermediate variables, the intermediate variable , and will be stored in shared memory. Then, each thread in the group will compute and store into registers only a fraction of the integrals. For example, the shape of the integral (dd—dd) is (6,6,6,6), each thread computes a fragment with the size of (3,3,3,3). 16 threads are needed to store the entire (dd—dd) integral. In the reduction stage, each thread is responsible for a fragment of 3x3 and . The fragments and are further reduced at the thread block level. The reduction scheme is presented in Algorithm 1. The shape of the intermediate variables are also known at compilation time. A similar idea has been proposed in (AAF+, 10), where is partitioned in , , and indices. is partitioned in and indices.
@@ 4.1. Compilation Overhead
| Integral pattern | Algorithm |
| (ss—ss)/(11—11) | 1Q1T | 284.8 | 25 | ||||
| (ss—ss)/(33—33) | 1Q1T | 512.0 | 74 | ||||
| (pp—pp)/(11—11) | 1Q1T | 638.8 | 75 | ||||
| (pp—pp)/(33—33) | 1Q1T | 880.6 | 94 | ||||
| (dd—dd)/(11—11) | 1QnT | 6011 | 194 | ||||
| (ff—ff)/(11—11) | 1QnT | 7697 | 270 | ||||
@@ 4.2. Benchmark with GPU4PySCF and TeraChem
| Molecule | Basis |
| gly30 | 6-31G* | 0.23 mHa | 0.0005 mHa | -0.07 mHa | ||||||
| gly30 | def2-TZVPP | 1.95 mHa | 0.0004 mHa | 2.10 mHa | ||||||
| valinomycin | 6-31G* | -0.06 mHa | -0.001 mHa | -0.26 mHa | ||||||
@@ 6. Future Work and Limitations
Author Contributions
Xiaojie Wu led the conceptualization of JIT algorithms and developed xQC v0.1, and wrote the initial draft. Qiming Sun contributed the improvements of the JK kernels v1.0 to v1.4 in GPU4PySCF, which are not documented elsewhere. Yuanheng Wang explored single-precision algorithms in GPU4PySCF, discussed JIT algorithms in xQC v0.1, and wrote the manuscript.
@@ Acknowledgements.
The authors appreciate the support of Dr. Wen Yan and the ByteDance Seed management team. Special thanks are also extended to Dr. Zhichen Pu for his discussions. The authors also thank the PySCF community for providing valuable tools that contributed to this work, especially the GPU4PySCF contributors.References
@@
- AAF+ [10] Andrey Asadchev, Veerendra Allada, Jacob Felder, Brett M. Bode, Mark S. Gordon, and Theresa L. Windus. Uncontracted rys quadrature implementation of up to g functions on graphical processing units. Journal of Chemical Theory and Computation, 6(3):696–704, 2010. PMID: 26613300.
@@
- DRK [76] Michel Dupuis, John Rys, and Harry F. King. Evaluation of molecular integrals over gaussian basis functions. The Journal of Chemical Physics, 65(1):111–116, 07 1976.